# Elasticity without a reference state: continuum mechanics of active tension nets

### v2

@@ Results
Figure 3: Isothermal coordinates “flatten” the tension manifold without shear. We isothermally map to the “natural domain” defined by the boundary condition $\lambda_{g}|_{\partial\Omega}=1$. The quasi-conformal map $w(z,\bar{z})$ maps the reference configuration to the physical domain. It determines the macroscopic stress tensor $\bm{\sigma}$ and is subject to physical boundary conditions (gray shading illustrates a clamped boundary, red arrows illustrate a traction force along the boundary).
@@
the macroscopic stress can be made to vanish identically: $\sigma_{\alpha\beta}^{\text{tot}}=\sigma_{\alpha\beta}-p\delta_{\alpha\beta}=0$. The $\mathbf{g}$-isothermal configuration, thus, is a stress-free reference configuration which emerges from a stress-only starting point. A liquid bubble raft provides a simple physical example: Surface tension balances against pressure so that the raft as a whole is stress-free (cutting it does not lead to macroscopic recoil). Equations (16) and (12) together imply a Laplace equation for the pressure in the stress-free configuration: $\Delta_{z}\log(p/p_{0})=-K_{g}$.
In a physical configuration, the pressure field $p$ is not an independent variable, but must obey the equation of state $p=P(n)$, Eq. (3). The stress-free reference state is physically attainable only if $p_{0}\lambda_{g}(\mathbf{z})=P(n(\mathbf{z}))$ where $n(\mathbf{z})$. For a flat tension metric, this relation reduces to
$n(\mathbf{z})=n({\bm{\xi}})\det[\partial_{\bm{\xi}}\mathbf{z}]\stackrel{{\scriptstyle?}}{{=}}n_{0}$ (17)
Eq. (17) does not hold in general. If it is not fulfilled, the $\mathbf{z}$-configuration is not mechanically balanced, and the tissue adopts a different configuration, in general, with residual stresses  [^2] Residual stresses are those present even for traction-free BCs. Note that the “incompatibility” of the tension manifold with the cell density that gives rise to residual stresses is distinct from the geometric incompatibility of a curved “reference metric” with a flat physical configuration found in the metric elasticity framework [52]. In the latter, the reference metric encodes local rest lengths in the material.. Nonetheless, the isothermal $\mathbf{z}$-configuration will be a useful reference to understand the resulting general force-balanced states. In particular, it will be the starting point for linearizing the theory.
@@
To relate the stress-free reference state $\mathbf{z}({\bm{\xi}})$ to the realized physical state, we need to calculate how the macroscopic tensile stress $\bm{\sigma}$ changes when the system is deformed from the isothermal reference into a new configuration, $z\mapsto w(z,\bar{z})$. Physically, a deformation moves and reorients the local active force dipoles (cell interfaces), thereby modifying the macroscopic stress tensor. Mathematically, $\mathbf{g}$ transforms according to Eq. (1). Using $g_{ab}(\mathbf{z})=\lambda_{g}(\mathbf{z})\delta_{ab}$, Eq. (4) becomes:
@@
The stress is thus determined by the square root of the Jacobian $\partial z^{\alpha}/\partial{w_{\beta}}$. In complexified notation:
$\begin{bmatrix}\sigma_{{\bar{w}}w}&\sigma_{{\bar{w}}{\bar{w}}}\\ \sigma_{ww}&\sigma_{w{\bar{w}}}\end{bmatrix}=2p_{0}\lambda_{g}\epsilon^{T}[\mathbf{D}w^{\dagger}\mathbf{D}w]^{-\frac{1}{2}}\epsilon$ (18)
@@
Eq. (20) relates spatial deformation to the changes in physical stress, and can be interpreted as an emergent stress-strain relationship. The anisotropic strain relative to the isothermal reference state sets the deviatoric stress $\sigma_{ww}$. Crucially, the stress anisotropy is not identical to the “microscopic” tension anisotropy (differences in interface tensions $\tau_{ij}$, or, geometrically, anisotropy of tension triangles). We will return to this important distinction in Sect. .4 where we define a tension anisotropy field in the continuum.
@@
$\bm{\sigma}$ is therefore symmetric and positive semi-definite, as required for a contractile/tensile stress.
However, not all deformations can lead to a mechanically balanced state. In complexified notation, force balance reads
$\bar{\partial}_{w}\sigma_{{\bar{w}}w}+\partial_{w}\sigma_{ww}$ $=2\bar{\partial}_{w}p$ (22)
As $\partial_{\alpha}\sigma_{\alpha\beta}=\partial_{\beta}p$, the net tensile force $\partial_{\alpha}\sigma_{\alpha\beta}$ must be curl-free:
$\partial_{w}\bar{\partial}_{w}\sigma_{{\bar{w}}w}+\partial_{w}^{2}\sigma_{ww}$ $=2\partial_{w}\bar{\partial}_{w}p$ (23)
The RHS is real, which entails a solvability condition:
@@
For a flat tension metric ($K_{g}=0$), one has $\lambda_{g}=1$, so this condition can be further simplified. In Cartesian coordinates (further assuming small displacement $|w-z|\ll 1$):
@@
That is, the curl of the deformation $w$ must be harmonic. Equation (24) constrains force-balanced physical configurations $w(z,\bar{z})$. Two classes of solutions are readily apparent.
@@
First, conformal maps $z=f(w)$, for which $\partial_{w}\bar{z}=\overline{\partial_{\bar{w}}z}=0$, so Eq. (24) holds trivially. Under conformal maps, $\sigma$ remains isotropic, while the pressure required for force balance transforms like $p\mapsto p/\lambda_{z|w}$. This allows broadening the conditions under which a stress-free configuration exists. If the reference state $\mathbf{z}({\bm{\xi}})$ fails to respect $n(\mathbf{z})=n_{0}$ (Eq. (17)), we can find a conformal map $w(z)$ so that $n(\mathbf{w})=\lambda_{z|w}^{-2}n(\mathbf{z})=n_{0}$, as long as $\Delta\log n(\mathbf{z})=0$ (which restriction is due to $\Delta\log\lambda_{z|w}=0$ for conformal maps).
As an aside, we note that the relation $p=p_{0}/\lambda_{z|w}$ between conformal deformations, which curve line elements, and pressure gradients can be understood as a continuum manifestation of the Young–Laplace law [13]. Consider a straight line with normal $\mathbf{n}$ in the reference configuration ($z$-coordinates). Through a conformal deformation it acquires a curvature $1/R=\mathbf{n}\cdot(2\bar{\partial}_{w}|\partial_{w}z|)$ (Appendix A), while the pressure drop $dp$ across the line element changes to $p_{0}\mathbf{n}\cdot(2\bar{\partial}_{w}|\partial_{w}z|)dr$. Recall that in $\mathbf{g}$-isothermal coordinates, $p_{0}dr=d\tau$ is the effective line tension, so that $d\tau/R=dp$ – a differential form of the Young–Laplace law.
@@
$\sigma_{w\bar{w}}=-8p_{0}\,\partial_{w}\bar{\partial}_{w}\theta,\;\sigma_{ww}=8p_{0}\,\bar{\partial}_{w}^{2}\theta$ (27)
An analogous, if lengthy, calculation shows that for $K_{g}\neq 0$, $z=\lambda_{g}^{-2}\bar{\partial}_{w}\theta$ is a perturbative solution to Eq. (24) (it corresponds to the Riemannian gradient of $\theta$ with respect to $\mathbf{g}$). The manifold of force-balanced states parametrized by $\theta$ corresponds to the “isogonal” modes of the discrete ATN model [22, 11], and we will therefore refer to $\theta$ as the isogonal potential (this connection is detailed in the companion paper [15]). Conformal and isogonal displacements exhaust all solutions to Eq. (24) (Appendix B). Any physical configuration $w$ can hence be obtained from the reference $z$ by a composition of a potential map and a conformal map. The two modes suffice to accommodate arbitrary deformations imposed on the system boundary (Appendix C).
@@
Appendix G.1 provides an elementary example of this mapping. While force balance in the bulk imposes a conformal–isogonal structure on the deformation $w$, BCs and a specific constitutive relation for pressure are required to fully determine the physical configuration.
@@
Purely isogonal deformations of the reference state also do not require pressure gradients, since the tensile stress is balanced on its own:
$\hskip-3.99994pt\bar{\partial}_{w}\sigma_{w\bar{w}}+\partial_{w}\sigma_{ww}=8p_{0}\,[\bar{\partial}_{w}(\partial_{w}\bar{\partial}_{w}\theta)-\partial_{w}(\bar{\partial}_{w}^{2}\theta)]=0$ (29)
Since the equation of state is trivial, $P(n)=p_{0}$, pressure also remains invariant. Isogonal modes in the bulk are therefore true soft modes (costing zero energy) in this case.
Under conformal deformations $w(z)$, however, the isotropic stress $\sigma_{w\bar{w}}=2p_{0}|\partial_{w}z|$ becomes non-uniform, implying pressure gradients. To sustain such pressure gradients, a finite compressibility (a non-constant $P(n)$) is required. Now, force balance $\partial_{\beta}p=\partial_{\alpha}\sigma_{\alpha\beta}$ selects a single physical solution among the isogonal–conformal family of solutions by fixing the isogonal potential $\theta$ in the bulk. Mechanical BCs further fix the boundary DOFs of the conformal and isogonal maps.
@@
Eq. (.2.4) is mathematically equivalent to linear isotropic elasticity [3] (even though the underlying physical picture is quite different). One identifies the effective shear and bulk moduli as $\frac{1}{2}p_{0}$ and $(B-1/2)p_{0}$. At higher orders in the gradient expansion, the elastic moduli scale like $\sim p$ and thus become non-uniform (high-pressure regions are stiffer). The finite shear modulus indicates that active tension networks are effective solids, rather than fluids. Furthermore, for a sufficiently large cellular bulk modulus $B$, both effective moduli are positive. The constitutive relation $P(n)$ thus ensures that the mechanical equilibrium is stable. If pressures are instead upgraded to independent, active degrees of freedom (like the tensions $\uptau_{ij}$), mechanical stability is no longer guaranteed. For example, fixing both pressure and surface tension in a soap bubble renders it unstable.
@@
As expected from the solvability analysis above, the general solution to Eq. (33) can be written as $u_{\alpha}=\partial_{\alpha}\theta+\nu_{\alpha}$ where $\nu_{\alpha}$ is a conformal vector field [23, 20]. Eq. (33) reduces to:
@@
where $C$ is an integration constant which can be absorbed into a uniform scaling. Using Eq. (34) and the Cauchy–Riemann equations for $\nu_{\alpha}$, the stress can be brought into the form:
@@
Gaussian curvature of the tension metric and $\delta n$ appear as source terms, implying residual stresses even for free BCs. These stresses result from an incompatibility between a constant cell density/pressure and an isotropic tensile stress (the macroscopically stress-free state does not have uniform pressure). The biharmonic equation Eq. (36), together with BCs on stress $\sigma_{\alpha\beta}$ or displacement $u_{\alpha}$, fully determines the tissue shape. It can be solved numerically or via complex variable techniques (referred to as Kolosov-Muskhelishvili formalism) [20]). Appendix G presents two worked examples, which illustrate the mapping from tension to physical space and the consequence of tension curvature.
A remarkable consequence of these results is that the shape and macroscopic mechanical properties of the tissue are determined by the tension metric and cell density alone. There are many microscopic tension configurations (tension triangulations) that have the same $\mathbf{g}$ and $n$. On a macroscopic level, they are all equivalent – the microscopic details are irrelevant for the static problem. Moreover, up to the overall scale $p_{0}$, the only free parameter of the theory is the cellular bulk modulus $B$.
@@
Let us briefly discuss four special cases: (i) For incompressible cells, $B\rightarrow\infty$ forcing $\partial_{\alpha}u_{\alpha}=-\delta n$. (ii) If pressures follow the ideal gas law $p(n)p_{0}n/n_{0}$, i.e. $B=1$, the total bulk and shear moduli are equal, which may be relevant to experiments on conventional foams with air-filled bubbles. (iii) For $B=1/2$, the total bulk modulus vanishes, so that the conformal mode $f_{\alpha}$ becomes soft. If $B<1/2$, the total bulk modulus is negative: intracellular pressure is insufficient to balance the isotropic tensile stress. Mechanical stability then requires a constraint on the total area. (iv) For fully compressible cells $B=0$, the isogonal potential $\theta$ is an unconstrained soft mode. However, conformal modes, which require pressure gradients, cannot be accommodated. By Eq. (36), freely compressible cells (i.e., $B=0$) are unstable if $K_{g}\neq 0$.
@@
In summary, boundary forces deform the tissue away from the emergent reference state $z(\xi,\bar{\xi})$. Incompatible cell densities and tension curvature lead to residual stresses, present even for free boundaries. The physical stress is governed by effective linear elasticity, leading to the biharmonic Eq. (36) for the isogonal potential, the effective Airy function. Due to Eq. (36), the physical configuration is the one with “minimal stress”, under the constraints imposed by BCs  [^4] Indeed, the biharmonic Eq. (36) can be written variationally: any solution must minimize the functional $\int[(\Delta\theta)^{2}+\theta\,\Delta(B^{-1}\log\lambda_{g}+\delta n)]d^{2}w$. Using $\mathrm{Tr}\bm{\sigma}^{\mathrm{tot}}=\Delta\theta$ and integrating by parts, the stress thus minimizes $\int[(\mathrm{Tr}\bm{\sigma})^{2}+(\mathrm{Tr}\bm{\sigma})(B^{-1}\log\lambda_{g}+\delta n)]d^{2}w$..
.3 Adiabatic dynamics: Beltrami flow
.3.1 Active stress dynamics
In summary, the map $\xi\mapsto w\big(\xi,\bar{\xi}\big)$ embeds the tension metric into physical space and defines the force-balanced state of the tissue. We next consider the temporal dynamics of the physical cell positions $w(\xi,{\bar{\xi}},t)$ in response to time-dependent active tensions. We assume rapid relaxation towards mechanical equilibrium. The tissue thus adopts the instantaneous equilibrium configuration $\mathbf{w}({\bm{\xi}})$, the solution of Eq. (5) (see App. D). We now consider it as a functional $w[\mathbf{g}(t)](\xi,\bar{\xi})$ of the time-dependent tension metric $\mathbf{g}(t)$. In this adiabatic approximation, the tissue flow due to tension dynamics can be computed by the chain rule:
$\partial_{t}w[\mathbf{g}(t)](\xi,{\bar{\xi}})$ $=\frac{\delta w[\mathbf{g}]}{\delta\mathbf{g}}\circ\partial_{t}\mathbf{g}(\xi,\bar{\xi})$
$:=\int\frac{\delta w[\mathbf{g}](\xi,\bar{\xi})}{\delta\mathbf{g}(\xi^{\prime},\bar{\xi}^{\prime})}\,\partial_{t}\mathbf{g}(\xi^{\prime},\bar{\xi}^{\prime})d\xi^{\prime}d\bar{\xi}^{\prime},$ (37)
@@
The functional derivative $\delta w/\delta\mathbf{g}$ is a convolutional integral operator, defined on the second line. This is a consequence of the non-local character of adiabatic flow, where a local change in contractility propagates across the tissue. We evaluate Eq. (37) by the composition of incremental maps and further decompose $w(\xi,\bar{\xi})$ into $\xi\mapsto z(\xi,\bar{\xi})\mapsto w(z,\bar{z})$, from $\xi$ to the reference $z$, and from $z$ to the physical state. By the chain rule:
@@
.3.2 Beltrami flow
@@
$\mathbf{g}({\bm{\xi}},t)\mapsto\mathbf{g}({\bm{\xi}},t+dt)=\mathbf{g}({\bm{\xi}},t)+\partial_{t}\mathbf{g}({\bm{\xi}},t)dt$ (39)
The resulting flow $z\mapsto z+\dot{z}\,dt$ must reestablish a state of isotropic stress. Expanding Eq. (10) in $dt$:
$\bar{\partial}_{\xi}\dot{z}(\xi,\bar{\xi},t)=\partial_{t}\mu_{g}(\xi,\bar{\xi},t)$ (40)

### v3

@@ Results
Figure 3: Isothermal coordinates “flatten” the tension manifold without shear. We isothermally map to the “natural domain” defined by the boundary condition $\lambda_{g}|_{\partial\Omega}=1$. The quasi-conformal map $w(z,\bar{z})$ maps the reference configuration to the physical domain. It determines the macroscopic stress tensor $\bm{\sigma}$ and is subject to physical boundary conditions (clamped boundary as gray shading, traction forces as red arrows).
@@
the macroscopic stress can be made to vanish identically: $\sigma_{\alpha\beta}^{\text{tot}}=\sigma_{\alpha\beta}-p\delta_{\alpha\beta}=0$. The $\mathbf{g}$-isothermal configuration, thus, is a stress-free reference configuration which emerges from a stress-only starting point. A liquid bubble raft provides a simple physical example (Appendix A). Surface tension balances against pressure so that the raft as a whole is stress-free (cutting it does not lead to macroscopic recoil). Equations (16) and (12) together imply a Laplace equation for the pressure in the stress-free configuration: $\Delta_{z}\log(p/p_{0})=-K_{g}$.
In a physical configuration, the pressure field $p$ is not an independent variable, but must obey the equation of state $p=P(n)$, Eq. (2). The stress-free reference state is physically attainable only if
$p_{0}\lambda_{g}(\mathbf{z})=P(n(\mathbf{z})).$ (17)
For a flat tension metric, this relation reduces to $n(\mathbf{z})=n({\bm{\xi}})\det[\partial_{\bm{\xi}}\mathbf{z}]=n_{0}$. If Eq. (17) is not fulfilled, the $\mathbf{z}$-configuration is not mechanically balanced. The tissue will adopt a different configuration with residual stresses that remain even when no traction is applied to the boundary  [^3] The “incompatibility” of the tension manifold with the cell density that gives rise to residual stresses is distinct from the geometric incompatibility of a curved “reference metric” with a flat physical configuration found in the metric elasticity framework [47]. In the latter, the reference metric encodes local rest lengths in the material.. Nonetheless, the isothermal $\mathbf{z}$-configuration will be a useful reference for the resulting general force-balanced states. In particular, it will be the starting point for linearizing the theory.
@@
To relate the stress-free reference state $\mathbf{z}({\bm{\xi}})$ to the realized physical state, we need to calculate how the macroscopic tensile stress $\bm{\sigma}$ changes when the system is deformed from the isothermal reference into a new configuration, $z\mapsto w(z,\bar{z})$. Physically, a deformation moves and reorients the local active force dipoles (cell interfaces), thereby modifying the macroscopic stress tensor. Mathematically, $\mathbf{g}$ transforms according to Eq. (1). Using $g_{ab}(\mathbf{z})=\lambda_{g}^{2}(\mathbf{z})\delta_{ab}$, Eq. (4) becomes:
@@
The stress is thus determined by the square root of the Jacobian $\partial z_{\alpha}/\partial{w_{\beta}}$. In complexified notation:
$\begin{bmatrix}\sigma_{{\bar{w}}w}&\sigma_{{\bar{w}}{\bar{w}}}\\ \sigma_{ww}&\sigma_{w{\bar{w}}}\end{bmatrix}=2p_{0}\lambda_{g}\bm{\epsilon}^{T}\cdot[\mathbf{D}w^{\dagger}\cdot\mathbf{D}w]^{-\frac{1}{2}}\cdot\bm{\epsilon}$ (18)
@@
Eq. (20) relates spatial deformation to the changes in physical stress, and can be interpreted as an emergent stress-strain relationship [22]. The anisotropic strain relative to the isothermal reference state sets the deviatoric stress $\sigma_{ww}$. Crucially, the stress anisotropy is not identical to the “microscopic” tension anisotropy (differences in interface tensions $\tau_{ij}$, or, geometrically, anisotropy of tension triangles). We will return to this important distinction in Sect. .4 where we define a tension anisotropy field in the continuum.
@@
In the companion paper [18], we derive the stress-metric relation in the form of Eq. (21) by explicit coarse graining.
In mechanical balance, the divergence of the tensile stress must be compensated by a pressure gradient. In complexified notation, force balance Eq. (5) reads
$\bar{\partial}_{w}\sigma_{{\bar{w}}w}+\partial_{w}\sigma_{ww}$ $=2\bar{\partial}_{w}p$ (22a)
$\Rightarrow\quad\partial_{w}\bar{\partial}_{w}\sigma_{{\bar{w}}w}+\partial_{w}^{2}\sigma_{ww}$ $=2\partial_{w}\bar{\partial}_{w}p$ (22b)
The RHS is real – the net tensile force must be curl-free. This entails a solvability condition:
@@
For a flat tension metric ($K_{g}=0$, so $\lambda_{g}=1$) and small displacement $|w-z|\ll 1$), this condition can be simplified. In Cartesian coordinates:
@@
The curl of the deformation $w$ must be harmonic. Equation (23) constrains force-balanced physical configurations $w(z,\bar{z})$. Two classes of solutions are readily apparent.
@@
First, conformal maps $z=f(w)$, for which $\partial_{w}\bar{z}=\overline{\partial_{\bar{w}}z}=0$, so Eq. (23) holds trivially. Under conformal maps, $\sigma$ remains isotropic, while the pressure required for force balance is $p\mapsto p/\lambda_{z|w}$. This allows broadening the conditions under which a stress-free configuration exists. If the reference state $\mathbf{z}({\bm{\xi}})$ fails to respect $n(\mathbf{z})=n_{0}$ (Eq. (17)), we can find a conformal map $w(z)$ so that $n(\mathbf{w})=\lambda_{z|w}^{-2}n(\mathbf{z})=n_{0}$, provided $\Delta\log n(\mathbf{z})=0$ (the restriction is due to $\Delta\log\lambda_{z|w}=0$).
As an aside, we note that the relation $p=p_{0}/\lambda_{z|w}$ between conformal deformations, which curve line elements, and pressure gradients can be understood as a continuum manifestation of the Young–Laplace law [12]. Consider a straight line with normal $\mathbf{n}$ in the reference configuration ($z$-coordinates). Through a conformal deformation it acquires a curvature $1/R=\mathbf{n}\cdot(2\bar{\partial}_{w}|\partial_{w}z|)$ (Appendix C), while the pressure drop $dp$ across the line element changes to $p_{0}\mathbf{n}\cdot(2\bar{\partial}_{w}|\partial_{w}z|)dr$. Recall that in $\mathbf{g}$-isothermal coordinates, $p_{0}dr=d\tau$ is the effective line tension, so that $d\tau/R=dp$ – a differential form of the Young–Laplace law.
@@
$\sigma_{w\bar{w}}=-2p_{0}\,\partial_{w}\bar{\partial}_{w}\theta,\;\sigma_{ww}=2p_{0}\,\bar{\partial}_{w}^{2}\theta$ (26)
An analogous, if lengthy, calculation shows that for $K_{g}\neq 0$, $z=\lambda_{g}^{-2}\bar{\partial}_{w}\theta$ is a perturbative solution to Eq. (23) (it corresponds to the Riemannian gradient of $\theta$ with respect to $\mathbf{g}$). The manifold of force-balanced states parametrized by $\theta$ corresponds to the isogonal “breathing” modes of the discrete ATN model [10], and we will therefore refer to $\theta$ as the isogonal potential (in the mathematical literature, such a potential defines a “weighted triangulation” [28]). Conformal displacements also have an exact equivalent at the cell-level – a certain “local” Möbius symmetry of force-balanced tilings. The discrete-continuum connections are detailed in the companion paper [18]. Conformal and isogonal displacements exhaust all solutions to Eq. (23) (Appendix D). Any physical configuration $w$ can hence be obtained from the reference $z$ by composing a potential map and a conformal map. The two modes suffice to accommodate arbitrary deformations imposed on the system boundary (Appendix E).
@@
Appendix B.1 provides an elementary example of this mapping. While force balance in the bulk imposes a conformal–isogonal structure on the deformation $w$, BCs and a specific constitutive relation for pressure are required to determine the physical configuration completely.
@@
Purely isogonal deformations of the reference state do not require pressure gradients, since the tensile stress is balanced on its own:
$\hskip-3.99994pt\bar{\partial}_{w}\sigma_{w\bar{w}}+\partial_{w}\sigma_{ww}=2p_{0}\,[\bar{\partial}_{w}(\partial_{w}\bar{\partial}_{w}\theta)-\partial_{w}(\bar{\partial}_{w}^{2}\theta)]=0$ (28)
Since the equation of state is trivial, $P(n)=p_{0}$, pressure also remains invariant. Isogonal modes in the bulk are therefore true zero-energy modes.
Under conformal deformations $w(z)$, however, the tensile stress $\sigma_{w\bar{w}}=2p_{0}|\partial_{w}z|$ becomes non-uniform, implying pressure gradients. To sustain such pressure gradients, a finite compressibility (a non-constant $P(n)$) is required. Now, force balance $\partial_{\beta}p=\partial_{\alpha}\sigma_{\alpha\beta}$ selects a single physical solution among the isogonal–conformal family of solutions by fixing the isogonal potential $\theta$ in the bulk. Mechanical BCs further fix the boundary DOFs of the conformal and isogonal maps.
@@
Existence and stability of force-balanced states.
Eq. (.2.4) is mathematically equivalent to linear isotropic elasticity [3] with an isotropic “body force” $p_{0}\partial_{\beta}(\log\lambda_{g}+B\delta n)$. However, the underlying physical picture is quite different: a conventional elastic material is akin to a Hookean spring, where strain relative to the rest length triggers a restoring force. In an active solid, active force dipoles, whose tension is independent of elastic strain, must be arranged (embedded) to balance against the isotropic pressure. The map to linear elasticity shows that a force-balanced embedding exists for any active tension configuration $\mathbf{g}$. This result relies on the existence of isothermal coordinates $z(\xi,\bar{\xi})$, in which the tensile stress is isotropic, and of two compatible modes (isogonal and conformal), which can accommodate elastic deformations in response to external and body forces.
One identifies the effective shear and bulk moduli as $\frac{1}{2}p_{0}$ and $(B-1/2)p_{0}$. At higher orders in the gradient expansion, the elastic moduli scale like $\sim p$ and thus become non-uniform (high-pressure regions are stiffer). The finite shear modulus indicates that active tension networks are effective solids, rather than fluids. Furthermore, for a sufficiently large cellular bulk modulus $B$, both effective moduli are positive. The constitutive relation $P(n)$ thus ensures that the mechanical equilibrium is stable. If pressures are instead upgraded to independent, active degrees of freedom (like the tensions $\uptau_{ij}$), mechanical stability is no longer guaranteed. For example, fixing both pressure and surface tension in a soap bubble renders it unstable.
Biharmonic equation and effective Airy function.
@@
As expected from the solvability analysis above, the general solution to Eq. (32) can be written as $u_{\alpha}=\partial_{\alpha}\theta+\nu_{\alpha}$ where $\nu_{\alpha}$ is a conformal vector field [29, 26]. Equation (32) reduces to:
@@
where $C$ is an integration constant which can be absorbed into a uniform scaling. Using Eq. (33) and the Cauchy–Riemann equations for $\nu_{\alpha}$, the stress reads:
@@
Gaussian curvature of the tension metric and $\delta n$ appear as source terms, implying residual stresses even for traction-free BCs. These stresses result from an incompatibility between a constant cell density and an isotropic tensile stress (the macroscopically stress-free state does not have uniform pressure). The biharmonic equation Eq. (35), together with BCs on stress $\sigma_{\alpha\beta}$ or displacement $u_{\alpha}$, fully determines the tissue shape. It can be solved numerically or via complex variable techniques (referred to as Kolosov-Muskhelishvili formalism) [26]). Appendix B presents two worked examples.
@@
Up to the overall scale $p_{0}$, the only free parameter of the theory is the cellular bulk modulus $B$. Let us briefly discuss four special cases: (i) For incompressible cells, $B\rightarrow\infty$ forcing $\partial_{\alpha}u_{\alpha}=-\delta n$. (ii) If pressures follow the ideal gas law $p(n)p_{0}n/n_{0}$, i.e. $B=1$, the total bulk and shear moduli are equal, which may be relevant to experiments on conventional foams with air-filled bubbles. (iii) For $B=1/2$, the total bulk modulus vanishes, so that the conformal mode $f_{\alpha}$ becomes soft. If $B<1/2$, the total bulk modulus is negative: intracellular pressure is insufficient to balance the isotropic tensile stress. Mechanical stability then requires a constraint on the total area. (iv) For fully compressible cells $B=0$, the isogonal potential $\theta$ is an unconstrained soft mode. However, conformal modes, which require pressure gradients, cannot be accommodated. By Eq. (35), freely compressible cells (i.e., $B=0$) are unstable if $K_{g}\neq 0$.
@@
In summary, boundary forces deform the tissue away from the emergent reference state $z(\xi,\bar{\xi})$. Incompatible cell densities and tension curvature lead to residual stresses, present even for free boundaries. The physical stress is governed by effective linear elasticity. The biharmonic Eq. (35) for the isogonal potential (the effective Airy function) implies that the physical configuration minimizes the physical stress, under the constraints imposed by BCs  [^5] Indeed, the biharmonic Eq. (35) can be written variationally: any solution must minimize the functional $\int[(\Delta\theta)^{2}+\theta\,\Delta(B^{-1}\log\lambda_{g}+\delta n)]d^{2}w$. Using $\mathrm{Tr}\bm{\sigma}^{\mathrm{tot}}=\Delta\theta$ and integrating by parts, the stress thus minimizes $\int[(\mathrm{Tr}\bm{\sigma})^{2}+(\mathrm{Tr}\bm{\sigma})(B^{-1}\log\lambda_{g}+\delta n)]d^{2}w$.. A remarkable consequence of these results is that the shape and macroscopic mechanical properties of the tissue are determined by the tension metric and cell density alone. There are many microscopic tension configurations (tension triangulations) that have the same $\mathbf{g}$ and $n$. On a macroscopic level, they are all equivalent – the microscopic details are irrelevant for the static problem.
.3 Active tension dynamics: Quasiconformal flow
We have seen how the cell positions $\xi\mapsto w\big(\xi,\bar{\xi}\big)$ embed the tension metric into physical space, and how they are determined by force balance. We next consider the dynamics of the embedding $w(\xi,{\bar{\xi}},t)$ in response to changes in active tension $\partial_{t}\mathbf{g}$. These dynamics are subject to biological control, which involves genetic programs and mechanochemical feedback loops [32, 33, 34, 35]. In the following, we do not assume any specific form, and instead consider a general active tension dynamics defined by
$\partial_{t}\log\mathbf{g}(\bm{\xi},t)=\mathbf{m}(\bm{\xi},t,\bm{\sigma},\cdots)$ (36)
The rate-of-change tensor $\mathbf{m}$ will act as an effective forcing term that internally drives the morphing of the active solid from one shape to another.
.3.1 Adiabatic approximation
We will assume rapid relaxation towards mechanical equilibrium. Consider relaxation dynamics of the cell positions $\mathbf{r}$
where $\Gamma$ is a friction coefficient. The stress $\sigma$ depends both on $\mathbf{r}(t)$ and $\mathbf{g}(t)$ by Eq. (4), and is thus time-dependent. The timescale of relaxation to equilibrium is given by $L\Gamma/p_{0}$, where $L$ is the system length scale. In the limit where relaxation is rapid (compared to tension dynamics $\partial_{t}\log\mathbf{g}$), Eq. (37) reduces to force balance:
$\partial_{\alpha}\sigma_{\alpha\beta}(t)-\partial_{\beta}p(t)\approx 0$ (38)
Due to rapid relaxation, transient friction or viscous stress will play no important role. The cells, therefore, adopt the instantaneous equilibrium configuration $\mathbf{w}[\mathbf{g}(t)](\xi,\bar{\xi})$, considered as a functional of the time-dependent tension metric $\mathbf{g}(t)$.
In the adiabatic approximation, the tissue flow due to tension dynamics can be computed by the chain rule:
$\partial_{t}w[\mathbf{g}(t)](\xi,{\bar{\xi}})$ $=\frac{\delta w[\mathbf{g}]}{\delta\mathbf{g}}\circ\partial_{t}\mathbf{g}=\frac{\delta w[\mathbf{g}]}{\delta\mathbf{g}}\circ\mathbf{m}(\xi,\bar{\xi})$
$:=\int\frac{\delta w[\mathbf{g}](\xi,\bar{\xi})}{\delta\mathbf{g}(\xi^{\prime},\bar{\xi}^{\prime})}\,\mathbf{m}(\xi^{\prime},\bar{\xi}^{\prime})d\xi^{\prime}d\bar{\xi}^{\prime}$ (39)
@@
The functional derivative $\delta w/\delta\mathbf{g}$ is a convolutional integral, defined on the second line. The adiabatic flow is thus “non-local”: a local change in contractility induces strain across the tissue. We evaluate Eq. (39) by the composition of incremental maps and further decompose $w(\xi,\bar{\xi})$ into $\xi\mapsto z(\xi,\bar{\xi})\mapsto w(z,\bar{z})$, from $\xi$ to the reference $z$, and from $z$ to the physical state. By the chain rule:
@@
.3.2 Adiabatic flow of the reference configuration
@@
$\mathbf{g}({\bm{\xi}},t)\mapsto\mathbf{g}({\bm{\xi}},t)+(\partial_{t}\mathbf{g}({\bm{\xi}},t))dt$ (41)
The resulting flow $z\mapsto z+\dot{z}\,dt$ must reestablish a state of isotropic stress. Rather than in Lagrangian $\xi$-coordinates, we express the adiabatic flow in the current $z(t)$ coordinates, so that $\dot{z}\bigr(z(t),\overline{z(t)},t\bigl)$ is the Eulerian velocity field. For brevity, we write $z=z(t)$. Expanding the Beltrami Eq. (10) in $dt$, and using the definition Eq. (36) of the $\mathbf{m}$-tensor leads to the simple expression
$\bar{\partial}_{z}\dot{z}(z,\bar{z})=\frac{1}{4}m_{zz}(z,\bar{z},t),\;\;\operatorname{Re}[\bar{\partial}_{z}\dot{z}]|_{\partial\Omega}=m_{z\bar{z}}|_{\partial\Omega}$ (42)